or a process with a mean of 100, a standard deviation of 10 and an upper specification of 120, what is the probability that a randomly selected item is defective(or beyond the upper specification limit)?


1
Expert's answer
2021-03-08T19:14:53-0500

"X\sim N(\mu, \sigma^2)"

Then "Z=\dfrac{X-\mu}{\sigma/}\sim N(0, 1)"

Given "\mu=100, \sigma=10"

"P(X>120)=1-P(X\leq 120)"

"=1-P(Z\leq\dfrac{120-100}{10})=1-P(Z\leq2)"

"\approx1-0.97725\approx0.02275"



"\dfrac{120-100}{10}=2"

95% of data fall with 2 standard deviations of the mean

"P(X>120)\approx\dfrac{1-0.95}{2}=0.025"